OpenStudy (anonymous):

The perimeter of an athletic field is 104 meters and the length is 18 m more than the width. What is the length and width

OpenStudy (zpupster):

Perimeter = w+w+l+l = 104 w= width and l = length but length = w+18 substitute this for length so now we have w+w+w+18+w+18 = 104 now add the w's together and the 18's also

OpenStudy (nurali):

Let x=length and y=width 2x+2y=104 x-18=y Substitute the 2nd equation into the first one 2x+2y=104 2x+2(x-18)=104 2x+2x-36=104 4x=140 x=140/4 x=35 Substitute x into one of the original equations x-18=y 35-18=y 17=y the length is 35m, the width is 17m

OpenStudy (nurali):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)